OpenStudy (agentnao):

A boat takes 2 hours longer to go up a 45 mi. long river than down it. If the boat's speed in still water is 12 mph, what is the current of the water? I think I know how to do this, but I can't fully remember. Can someone nudge me in the right direction?

OpenStudy (anonymous):

asuume velcity of current to be a constant x find the differrence in time taken by boat in current and in still water to cover 45mi and equate eith given differrence in time and solve for x

OpenStudy (agentnao):

So, in still water, the boat takes 3.75 hours to travel 45 miles. And I would then subtract that from 45x?

OpenStudy (agentnao):

Should I be using the d=rt formula? Or something else?

OpenStudy (dumbcow):

distance = rate*time Up river, c is speed of current 45 = (12-c)*t Down river 45 = (12+c)(t-2)

OpenStudy (agentnao):

I could set those equal to each other, but there are two variables. How would I solve for both? Also, wouldn't it be t+2 going down river because it takes 2 hours longer?

OpenStudy (dumbcow):

no it takes 2 hours longer to go up river against the current you could do substitution here t = 45/(12-c)

OpenStudy (anonymous):

Sorry for mistake in my answer . i thought that the case dealt with boat in still water and downstream not upstream and downstream. the time taken by boat upstream =45/12-x the time taken by boat downstream is = 45/12+x now find difference and equate with 2 solve x

OpenStudy (agentnao):

Oh, oops! Sorry, I read that wrong. It is t-c. And hm. So I could substitute that into the other equation, solve for c and then replug that value into one of the equations to find t. Okay, I'm going to try it right now and let you know what I get.

OpenStudy (dumbcow):

ok

OpenStudy (agentnao):

I don't think the substituting t thing works.

OpenStudy (dumbcow):

it does but its a little messy i guess you get stuck try it vicky's way

OpenStudy (agentnao):

Yes, definitely a bit messy. Alright, I'll try hers too.

OpenStudy (anonymous):

(45/12-x) - (45/12+x)=2 45x=288-x^2 then x=3 or -48

OpenStudy (anonymous):

you know what i am not a girl

OpenStudy (agentnao):

*his, my apologies

OpenStudy (dumbcow):

haha

OpenStudy (agentnao):

How do you know it equals two?

OpenStudy (anonymous):

its the difference in time

OpenStudy (agentnao):

Hm. So the difference between the rate with the current and the rate without the current equals 2. What is the 288 from?

OpenStudy (anonymous):

i am sorry its not 288 its 144

OpenStudy (agentnao):

Can you show me step by step in this whiteboard room?

OpenStudy (agentnao):

I think you lost me somewhere.

OpenStudy (agentnao):

http://www.twiddla.com/599175

OpenStudy (dumbcow):

\[45 = (12+c)(\frac{45}{12-c}-2)\] \[45 = \frac{45(12+c)}{12-c}-2(12+c)\] \[45 = \frac{45(12+c)}{12-c}-\frac{2(12+c)(12-c)}{12-c}\] \[45 = \frac{45(12+c)-2(12+c)(12-c)}{12-c}\] \[45(12-c) = 45(12+c) -2(12+c)(12-c)\] \[540-45c = (540+45c)-2(144-c^{2})\] \[540-45c = (540+45c)-288+2c^{2}\] \[2c^{2}+90c-288=0\] \[c^{2}+45c-144=0\] \[(c+48)(c-3) = 0\] \[c=3\]

OpenStudy (dumbcow):

hopefully you can find where you got stuck :)

OpenStudy (agentnao):

Wow, that is a wonderful step by step answer! Thank you very much :) I've looked over it just now, but I'll check it out in more detail tomorrow because it's very late over here. Thanks again though and if I have any other questions about this, I'll bother you again :)

OpenStudy (dumbcow):

no problem night

OpenStudy (agentnao):

dumbcow? Please please please tell me you're on right now?

OpenStudy (agentnao):

vicky? what about you?

OpenStudy (dumbcow):

yes?

OpenStudy (agentnao):

Oh, never mind. Ishaan helped me instead. Thank you though.

OpenStudy (dumbcow):

ok

OpenStudy (agentnao):

Actually, I take that back. Ishaan seems to be busy elsewhere, so I could still use your help. He answered the problem, but I still have some questions about it.

OpenStudy (dumbcow):

haha ok sorry i didn't notice your post, post the link to your question and i'll try to answer your questions